OpenStudy (anonymous):

I don't understand word problems :/ A copy machine makes 32 copies per minute. How many copies does it make in 5 minutes and 15 seconds?

OpenStudy (texaschic101):

I am changing the 5 min 15 sec to sec....(5)(60) + 15 = 315 sec now make a proportion 32/60 = x/315 (32 copies to 60 sec = x copies to 315 sec) 60x = 10080 x = 10080/60 x = 168 copies

OpenStudy (anonymous):

yes exactly

OpenStudy (anonymous):

Okay . . . . I Kinda Get It But I Kinda Don't. Like I Got The First Two Steps But How Did You Get 10080.? By 60*315.?

OpenStudy (texaschic101):

32/60 = x/315 cross multiply (60)(x) = (315)(32) understand now ?

OpenStudy (anonymous):

OH.! Okay Now I Understand How You Got That.!

OpenStudy (texaschic101):

good....in proportions, you cross multiply

OpenStudy (anonymous):

Okay Well Thank You.! Now I Fully Understood That. My Teacher Taught Us Completely Different And Very Complicated.